A certain radioactive nuclide has a half life of 74.8 years. How long does it take for 87.5% of a sample of this substance to decay?

___years?
Answer #1

In (SA) = kt a time t At initial concentration At concentration after k is rate constant K = 0.693 It is the half-life. InCARemaining percent = 100-87.5 12.5 % 100 A = 12.5 CAO CA = 0.125 CAO CAO 0-693 Los conoscere ) = 0612x+ 0.125 (Ao 74.8 year in